OpenStudy (anonymous):

can someone help me find the domain and range please??? (-1, 1/2), (-1/2, -1), (3/2, 0), (2, 3/2)

OpenStudy (anonymous):

domain: list the first numbers range: list the second numbers

OpenStudy (anonymous):

domain:\(\{-1,-\frac{1}{2},\frac{3}{2},2\}\) i will let you do the range on your own

OpenStudy (anonymous):

1/2 -1 0 3/2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

i didnt find the answer.. i have to graph the relation but i cant find the relation :((

OpenStudy (anonymous):

i don't know what you mean exactly you asked for the domain and range, not the graph to write the graph, you can only plot those four points

OpenStudy (anonymous):

i thought i would be able to graph that after finding the domain and range but i cant..

OpenStudy (anonymous):

i dont know how to graph it..

OpenStudy (anonymous):

just plot the four points

OpenStudy (anonymous):

\((-1\frac{1}{2})\)

OpenStudy (anonymous):



OpenStudy (anonymous):

easier with graph paper

OpenStudy (anonymous):

oh okay, thank you!!! again ^^

OpenStudy (anonymous):

yw